# Toward Robust, Reproducible, and Widely Accessible Intracranial Language Brain-Computer Interfaces: A Comprehensive Review of Neural Mechanisms, Hardware, Algorithms, Evaluation, Clinical Pathways and Future Directions

### 03 Mar 2026 v1

@@ 2.2 Hierarchical Functional Streams in Speech Processing
Intracranial recordings provide evidence for distributed networks encoding phonetic, spectrotemporal, and lexical-semantic representations [108, 125, 53]. For spectrotemporal processing, posteromedial Heschl’s gyrus (HG), the presumed core auditory cortex, resolves the temporal envelope of speech stimuli using mechanisms across a wide range of ECoG frequencies, with high-frequency event-related band power (ERBP) reflecting fine-grained temporal encoding even under compromised intelligibility [93]. High gamma power (70–150 Hz) tracking of speech is concentrated in the superior temporal gyrus (STG), while low-frequency phase tracking (1–7 Hz) is more widespread, including STG, anterior temporal cortex, and inferior frontal cortex [38]. Additionally, bilateral temporal and motor cortices encode speech envelopes, with the left STG specializing in fine-grained phonetic information and the right middle temporal gyrus (MTG) in slower, syllable-level representations [56]. Beyond cortical areas, deeper structures like the hippocampus and thalamus contribute to speech production, as shown by SEEG recordings covering diverse cortical and subcortical regions [121].
@@
Phonetic representations are encoded in distributed neural populations. The STG exhibits a caudal-anterior functional gradient, with posterior "onset" zones sensitive to acoustic onsets and phrase boundaries, and anterior "sustained" zones encoding ongoing phonetic and lexical-semantic features [43]. Within STG, phonetic feature selectivity is organized by manner and place of articulation, with distinct clusters for obstruents (plosives, fricatives) and sonorants (vowels, nasals), forming a multidimensional acoustic feature space [87]. During speech production, phonetic features can be decoded from neural activity in vSMC and STG, with confusions clustering by articulatory characteristics [90]. The face motor cortex (FMC) shows locally distributed neural assemblies encoding articulatory movements, enabling higher word classification accuracy compared to Wernicke’s area, which exhibits more abstract semantic encoding [62]. Phonetic discriminability during continuous speech production involves dynamic activation of Broca’s area (planning), sensorimotor areas, and STG (processing), with peak activation timing aligned to phone production [50]. Speech decoding leverages both articulatory (sensorimotor cortex) and perceptual/semantic (temporal lobe) information, with vSMC often providing the most informative signals, and increased electrode coverage improving accuracy [114].
@@
Lexical-semantic processing engages ventral stream regions. Word prediction primarily activates lower-order regions like STG and MTG, while sentence prediction involves default mode network (DMN) regions such as the temporoparietal junction (TPJ), medial prefrontal cortex (mPFC), and precuneus [141, 49]. Lexical activation is associated with left posterior MTG (pMTG), and lexical selection with left inferior frontal gyrus (LIFG) and medial frontal areas, with left pMTG lesions causing persistent word retrieval deficits [104]. Semantic information can be decoded from intracranial recordings in inferior frontal gyrus (IFG) and STG, supporting the ventral stream’s role in form-to-meaning mapping.
@@
Cross-stream interaction is evident in overlapping functional networks and dynamic information flow. The STG’s onset and sustained zones interact to support speech parsing, with onset activity providing temporal landmarks and sustained activity enabling continuous phonetic processing. Functional interactions between word and sentence prediction levels occur sparsely at sentence boundaries, supporting the sparse updating hypothesis. Conduction aphasia, resulting from impaired integration between dorsal and ventral streams, highlights the necessity of cross-stream communication for fluent speech. Recurrent neural computations within STG may enable binding of form to articulation and meaning, making stream reliance context-sensitive [136].
@@
Individual differences in stream localization are observed across spatial, spectral, and functional domains. In HG, envelope following shows considerable intersubject variability, with some subjects maintaining high-frequency ERBP envelope tracking even for unintelligible speech. Gamma reactivity frequencies (low vs high gamma) and spatial topography vary across subjects, reflecting individual differences in functional anatomy [26]. STG phonetic feature clusters show systematic individual variation but share common organizational principles. Speech-responsive electrode distributions, including tone- and syllable-discriminative electrodes in vSMC and widespread speech-responsive regions, exhibit inter-subject variance influenced by language behaviors and cortical localization [137]. Sensorimotor cortex activation patterns during speech perception overlap with production but show individual-specific differences, leading to potential false positives in language BCIs trained on production data [107].
@@
The signal gradient of different speech types modulates reliance on each stream. High gamma power, associated with bottom-up acoustic encoding (dorsal stream), peaks early (<100 ms), while low-frequency phase responses, linked to hierarchical integration (ventral stream) [36], have later peaks ( 150 ms) [38]. Phonological production tasks rely more on the dorsal stream, while lexical-semantic comprehension tasks depend on the ventral stream. Decoding latency and speed vary with speech type and vocabulary complexity, indicating modulation of stream utilization [74]. Different speech stimuli (phonemes, words, sentences) engage distinct hierarchically organized networks [42], with acoustic-phonetic features driving dorsal stream reliance and lexical-semantic features enhancing ventral stream engagement [127].
@@ 2.3 Neural Signatures of Overt, Mimed, and Imagined Speech
Overt, mimed, and imagined speech exhibit distinct neural correlates across frequency bands and cortical regions. For instance, power spectrum changes for overt and imagined speech are spatially comparable but less pronounced for imagined speech, with fewer cortical sites showing significant changes, particularly in broadband high-frequency activity (BHA). BHA increases in sensory and motor regions for both overt and imagined speech but only in the superior temporal cortex during overt speech, decreasing during imagined speech, likely due to the absence of auditory feedback [100, 29]. ECoG studies have demonstrated distinct spatio-temporal cortical activation patterns during continuous overt and covert speech production: fronto-motor regions (ventral primary motor cortex, ventral premotor cortex, inferior frontal gyrus) activate prior to speech output during overt speech, while temporal regions (superior and middle temporal gyri) activate primarily during and after speech output, reflecting auditory feedback processing [10, 63, 94]. Direct comparison of overt and covert speech ECoG signals reveals common activations in auditory cortex, premotor and primary motor cortex, frontal eye fields, and dorsolateral prefrontal cortex, but covert speech signals do not reliably encode acoustic features like the speech intensity envelope. Magnetoencephalography (MEG) studies further show distinct decoding performances across stages (pre-stimuli, perception, preparation/imagination, production/articulation) for overt and covert speech, with the highest accuracy during production, followed by imagination [28]. Additionally, speech imagery, including covert and imagined speech, involves distinct neural mechanisms engaging Broca’s and Wernicke’s areas with differential activation patterns [115, 123].
@@
Evidence of hierarchical channel nesting and differential frequency contributions to decoding is substantial. While BHA provides the best signal for overt speech decoding, both low- and higher-frequency power and local CFC contribute importantly to imagined speech decoding, particularly in phonetic and vocalic spaces. Low-frequency power (e.g., theta, low-beta) and cross-frequency dynamics (e.g., phase-amplitude coupling between low-beta and BHA) contain key information for imagined speech decoding. Spatial-spectral-temporal convolutional neural network (CNN) approaches, which integrate spatial information from multiple sensors with spectral and temporal features, enhance decoding accuracy by capturing hierarchical channel nesting and whole-brain dynamics. Intracranial electroencephalography (iEEG) studies identify distinct frequency bands: high-frequency gamma band activity linked to phoneme articulation and low-frequency theta band activity associated with syllabic and prosodic modulation, with cross-frequency coupling (e.g., theta phase-gamma amplitude coupling) reflecting hierarchical neural dynamics [2]. Thalamocortical studies further reveal hierarchical processing, where thalamic neurons prefer higher temporal modulation rates, and cortical neurons show slower preferences, indicating differential frequency contributions across regions [89]. For overt speech, gamma band (70-170 Hz) activity is highly task-related, with spectrogram reconstruction showing correlations peaking around the first formant frequency ( 300 Hz) [51]. Combining low-frequency components (0–50 Hz) with the high-gamma envelope (70–150 Hz) yields higher speech reconstruction accuracy than single bands, implicating complementary hierarchical encoding [1].
@@
Controversies persist regarding how far imagined-speech decoding can scale beyond constrained paradigms. Supportive evidence includes low-/cross-frequency intracranial decoding above chance [100], word-pair classification from direct cortical recordings [82], non-invasive phrase decoding during imagination stages [28], and recent real-time attempted/imagined-speech decoding in implanted users [66]. Counter-evidence includes weak acoustic-envelope recoverability in covert conditions [10], performance drops in earlier continuous speech pipelines [50], and systematic reviews emphasizing heterogeneity in tasks, compliance checks, and preprocessing choices [115]. Methodologically, limited paired neural-speech supervision and temporal misalignment remain practical bottlenecks [19, 121, 47, 139]. Overall, current studies indicate feasibility for proof-of-concept tasks but suggest that clinical-grade reliability remains to be established.
@@
Across multiple datasets, an SNR gradient is observed from overt to mimed (silently attempted without vocalization) to imagined speech, with overt conditions typically exhibiting stronger and more spatially extensive high-gamma (BHA) activity than imagined conditions [100, 10, 112]. This gradient both motivates and constrains imagined-speech decoding: positive feasibility evidence shows above-chance covert/imagined decoding in constrained word/phrase paradigms using low-frequency and cross-frequency-coupling (CFC) features [100, 82, 28], and recent implanted-participant work further reports real-time decoding of both attempted and imagined speech [66]. At the same time, negative findings indicate that covert signals often encode acoustic detail less reliably and vary substantially across participants and tasks [10, 50, 115], suggesting that robust large-vocabulary communication remains to be established. Crucially, the SNR gradient guides the choice of intermediate representations. Overt speech-related motor signals are stronger and more directly linked to acoustic features, supporting decoding strategies that exploit high-frequency articulatory correlates [81, 16]. In contrast, imagined speech exhibits weaker, more spatially diffuse activity and poorer discriminability of articulatory representations in BHA; instead, it more reliably engages perceptual representations (phonetic and vocalic) indexed by low-frequency bands and CFC [100, 75, 29], consistent with the flexible abstraction hypothesis [25]. Given the weaker SNR and typical data scarcity in implanted settings, this motivates targeting low-frequency and CFC features [100, 75] and decoding through explicit articulatory–acoustic intermediate parameters (e.g., pitch, formant frequencies) [8] rather than direct spectrograms or latent vectors, which can reduce sample complexity and improve trainability with limited neural data [19, 4]. Thus, imagined-speech variability necessitates feature- and representation-level adaptation toward perceptual/low-frequency structure [100, 25, 29], whereas overt speech can capitalize on higher-SNR, high-frequency articulatory signals [81, 16].
@@ 2.4 Temporal Dynamics and Context Integration in Speech Representation
Temporally recurrent connections within cortical regions play a crucial role in generating context-dependent phonological representations, a mechanism mirrored in computational models and neural recordings. In 1997, Hochreiter and Schmidhuber [52] introduced Long Short-Term Memory (LSTM) networks, which use memory cells with constant error carousels and multiplicative gate units to control information flow, enabling the storage of context-dependent information over long time intervals and supporting the generation of dynamic, context-sensitive phonological representations. Building on this, in 2006, Graves et al. [40] demonstrated that bidirectional LSTM (BiLSTM) networks, combined with Connectionist Temporal Classification (CTC) loss, implicitly model inter-label dependencies by predicting labels that commonly occur together in connected patterns, reflecting how cortical recurrent connections might integrate temporal context to shape phonological encoding. Empirically, in 2022, Luo et al. [79] highlighted that deep learning architectures like BiLSTMs map temporally structured neural features to sequence-based speech representations, with temporally recurrent neural connections within cortical regions providing the substrate for context-dependent integration in speech representation [92, 125]. This is further supported by 2023, Metzger et al. [88], who observed that the sensorimotor cortex (SMC) exhibits temporally recurrent articulatory representations somatotopically organized by place of articulation (POA), directly contributing to context-dependent phonological encoding.
@@ 3.1 Intracranial Electrode Families, Representative Suppliers, and Development Trajectories
The trajectory of intracranial electrodes in language BCI can be summarized as a progression from clinically available mapping hardware to purpose-optimized neuroprosthetic interfaces. Early speech-decoding studies largely repurposed epilepsy-monitoring electrodes (macro-ECoG grids/strips and later SEEG depth electrodes), enabling proof-of-concept decoding while leveraging established neurosurgical workflows and clinical monitoring infrastructure [70, 69, 50]. This stage established that useful speech information is present in cortical field potentials, especially in high-gamma activity, and that intracranial recordings can support continuous or near-continuous decoding under controlled tasks.
The next phase focused on improving spatial resolution, task complexity, and decoding fidelity. High-density cortical recordings and improved modeling pipelines enabled stronger articulation-level decoding and speech synthesis, while broader datasets clarified how electrode placement and coverage heterogeneity shape performance [5, 80, 91, 79, 33]. In parallel, SEEG expanded the accessible anatomical search space beyond cortical surface coverage, supporting distributed language-network sampling and more flexible trajectory planning under clinical constraints [19, 132, 99]. Recent work represents a third phase in which electrode hardware is increasingly evaluated by chronic utility rather than acute signal quality alone. Intracortical MEAs enabled high-rate and high-fidelity speech neuroprosthesis demonstrations [128, 12, 130], while chronic ECoG systems demonstrated stronger home-use feasibility in selected users with ALS [78, 4]. This shift reframes “best electrode” as a use-case-dependent question balancing coverage, resolution, packaging, calibration burden, and care logistics rather than maximizing any single signal metric [114, 110, 133].
@@
Intracortical MEAs have also shown significant advancements in capturing speech-related neural activity. In 2023, Willett et al. [128] used high-resolution spiking activity recordings from intracortical arrays to achieve a 9.1% WER on a 50-word vocabulary and 23.8% on a 125,000-word vocabulary, with decoding speeds of 62 words per minute. They further identified that the ventral premotor cortex (area 6v) contains a dense, intermixed representation of speech articulators within a small (3.2 × 3.2 mm²) region, with phoneme articulation details preserved even years after paralysis. Notably, decoding accuracy improved log-linearly with electrode count, suggesting that denser arrays could further enhance performance. Similarly, in 2025, Wood et al. [130] implanted 256 microelectrodes in the precentral gyrus, capturing neural activity that drove a voice synthesizer to generate speech resembling the participant’s pre-ALS voice with  25 ms latency, including paralinguistic features like intonation and emphasis for modulating speech style (e.g., statements vs. questions). For ECoG, in 2022, Luo et al. [79] highlighted innovations such as 32 recording contacts per electrode, robotic insertion techniques to minimize tissue trauma, and flexible array designs, which address traditional limitations in spatial density and coverage. These advances, combined with deep learning methods mapping ECoG high gamma activity to vocal tract trajectories, support improved neural capture for speech decoding. While long-term ECoG stability for speech remains under investigation, motor BCI research has demonstrated reliable chronic decoding, and safety in individuals with late-stage ALS has been reported.
These emerging technologies, including MoS2-based active arrays with ultrahigh spatiotemporal resolution, high-channel-count intracortical microelectrodes, and advanced ECoG configurations, overcome traditional constraints by enhancing resolution, biocompatibility, and coverage. They enable capture of fine-grained speech-related neural activity, from spiking dynamics to paralinguistic features, and hold significant potential for advancing clinical translation of language BCIs.

### 16 Mar 2026 v2

@@ 2.2 Hierarchical Functional Streams in Speech Processing
Intracranial recordings provide evidence for distributed networks encoding phonetic, spectrotemporal, and lexical-semantic representations [110, 127, 55]. For spectrotemporal processing, the posteromedial Heschl’s gyrus (HG), the presumed core auditory cortex, resolves the temporal envelope of speech stimuli through mechanisms across a wide range of ECoG frequencies. High-frequency event-related band power (ERBP) reflects fine-grained temporal encoding in this region, even under conditions of compromised intelligibility [95]. High gamma power (70–150 Hz) tracking of speech is concentrated in the superior temporal gyrus (STG), while low-frequency phase tracking (1–7 Hz) is more widespread, encompassing the STG, anterior temporal cortex, and inferior frontal cortex [39]. Bilateral temporal and motor cortices encode speech envelopes, with the left STG specializing in fine-grained phonetic information and the right middle temporal gyrus (MTG) in slower, syllable-level representations [58]. Beyond cortical areas, deeper structures like the hippocampus and thalamus contribute to speech production, as shown by SEEG recordings covering diverse cortical and subcortical regions [123].
@@
Phonetic representations are encoded in distributed neural populations. The STG exhibits a caudal-anterior functional gradient, with posterior "onset" zones sensitive to acoustic onsets and phrase boundaries, and anterior "sustained" zones encoding ongoing phonetic and lexical-semantic features [44]. Within STG, phonetic feature selectivity is organized by manner and place of articulation, with distinct clusters for obstruents (plosives, fricatives) and sonorants (vowels, nasals), forming a multidimensional acoustic feature space [89]. During speech production, phonetic features can be decoded from neural activity in the vSMC and STG, with confusion patterns clustering according to articulatory characteristics [92]. The face motor cortex (FMC) shows locally distributed neural assemblies encoding articulatory movements, enabling higher word classification accuracy than Wernicke’s area, which exhibits more abstract semantic encoding [64]. Phonetic discriminability during continuous speech production involves dynamic activation of Broca’s area (supporting articulation planning), sensorimotor areas (supporting articulation), and STG (supporting speech processing), with peak activation timing aligned to phone production [52]. Speech decoding leverages both articulatory information from the sensorimotor cortex and perceptual or semantic information from the temporal lobe. The vSMC often provides the most informative signals, with increased electrode coverage further improving decoding accuracy [116].
@@
Lexical-semantic processing engages ventral stream regions. Word prediction primarily activates lower-order regions like the STG and MTG, while sentence-level prediction recruits default mode network (DMN) regions, such as the temporoparietal junction (TPJ), medial prefrontal cortex (mPFC), and precuneus [143, 51]. Lexical activation is associated with the left posterior MTG (pMTG), while lexical selection engages the left inferior frontal gyrus (LIFG) and medial frontal areas. Lesions to the left pMTG result in persistent word retrieval deficits [106]. Semantic information can be decoded from intracranial recordings in the inferior frontal gyrus (IFG) and STG, supporting the ventral stream’s role in form-to-meaning mapping.
@@
Cross-stream interaction is evident in overlapping functional networks and dynamic information flow. The STG’s onset and sustained zones interact to support speech parsing, with onset activity providing temporal landmarks and sustained activity enabling continuous phonetic processing. Functional interactions between word and sentence prediction levels occur sparsely at sentence boundaries, supporting the sparse updating hypothesis. Conduction aphasia, arising from impaired integration between dorsal and ventral streams, underscores the critical role of cross-stream communication for fluent speech. Recurrent neural computations within the STG may enable the binding of form to articulation and meaning, rendering stream reliance context-sensitive [138].
@@
Individual differences in stream localization are observed across spatial, spectral, and functional domains. In HG, envelope following shows considerable intersubject variability, with some subjects maintaining high-frequency ERBP envelope tracking even for unintelligible speech. The frequency profile of gamma reactivity (low vs high gamma) and its spatial topography vary across subjects, reflecting individual differences in functional anatomy [27]. STG phonetic feature clusters show systematic individual variation but share common organizational principles. Speech-responsive electrode distributions, including tone- and syllable-discriminative electrodes in vSMC and widespread speech-responsive regions, exhibit intersubject variance influenced by language behaviors and functional anatomy [139]. Sensorimotor cortex activation patterns during speech perception overlap with those observed during production but show individual-specific differences, potentially leading to false-positive activations in language BCIs trained on production data [109].
@@
The signal gradient across different speech types modulates the degree of reliance on each stream. High gamma power, associated with bottom-up acoustic encoding in the dorsal stream, peaks early (<100 ms), while low-frequency phase responses, linked to hierarchical integration in the ventral stream [37], peak later ( 150 ms) [39]. Phonological production tasks rely more on the dorsal stream, while lexical-semantic comprehension tasks depend on the ventral stream. Decoding latency and speed vary with speech type and vocabulary complexity, indicating that stream utilization is dynamically modulated [76]. Different speech stimuli, such as phonemes, words, and sentences, engage distinct hierarchically organized networks [43], with acoustic-phonetic features driving dorsal stream reliance and lexical-semantic features enhancing ventral stream engagement [129].
@@ 2.3 Neural Signatures of Overt, Mimed, and Imagined Speech
Overt, mimed, and imagined speech exhibit distinct neural correlates across frequency bands and cortical regions. For instance, power spectrum changes for overt and imagined speech are spatially comparable but less pronounced for imagined speech, with fewer cortical sites showing significant changes, particularly in broadband high-frequency activity (BHA). BHA increases in sensory and motor regions for both overt and imagined speech, but only in the superior temporal cortex during overt speech, decreasing during imagined speech, likely due to the absence of auditory feedback [102, 30]. ECoG studies have demonstrated distinct spatio-temporal cortical activation patterns during continuous overt and covert speech production: fronto-motor regions (ventral primary motor cortex, ventral premotor cortex, inferior frontal gyrus) activate prior to speech output during overt speech, while temporal regions (superior and middle temporal gyri) activate primarily during and after speech output, reflecting auditory feedback processing [10, 65, 96]. Direct comparison of overt and covert speech ECoG signals reveals common activations in auditory cortex, premotor and primary motor cortex, frontal eye fields, and dorsolateral prefrontal cortex, but covert speech signals do not reliably encode acoustic features like the speech intensity envelope. Magnetoencephalography (MEG) studies further show distinct decoding performances across stages (pre-stimuli, perception, preparation/imagination, production/articulation) for overt and covert speech, with the highest accuracy during production, followed by imagination [29]. Additionally, speech imagery, including covert and imagined speech, involves distinct neural mechanisms engaging Broca’s and Wernicke’s areas with differential activation patterns [117, 125].
@@
Evidence of hierarchical channel nesting and differential frequency contributions to decoding is substantial. While BHA provides the best signal for overt speech decoding, both low- and higher-frequency power and local cross-frequency-coupling (CFC) contribute importantly to imagined speech decoding, particularly in phonetic and vocalic spaces. Low-frequency power (e.g., theta, low-beta) and cross-frequency dynamics (e.g., phase-amplitude coupling between low-beta and BHA) contain key information for imagined speech decoding. Spatial-spectral-temporal convolutional neural network (CNN) approaches, which integrate spatial information from multiple sensors with spectral and temporal features, enhance decoding accuracy by capturing hierarchical channel nesting and whole-brain dynamics. Intracranial electroencephalography (iEEG) studies identify distinct frequency bands: high-frequency gamma band activity linked to phoneme articulation and low-frequency theta band activity associated with syllabic and prosodic modulation, with cross-frequency coupling (e.g., theta phase-gamma amplitude coupling) reflecting hierarchical neural dynamics [2]. Thalamocortical studies further reveal hierarchical processing, where thalamic neurons prefer higher temporal modulation rates, and cortical neurons show slower preferences, indicating differential frequency contributions across regions [91]. For overt speech, gamma band (70-170 Hz) activity is highly task-related, with spectrogram reconstruction showing correlations peaking around the first formant frequency ( 300 Hz) [53]. Combining low-frequency components (0–50 Hz) with the high-gamma envelope (70–150 Hz) yields higher speech reconstruction accuracy than single bands, implicating complementary hierarchical encoding [1].
@@
Controversies persist regarding how far imagined-speech decoding can scale beyond constrained paradigms. Supportive evidence includes low-/cross-frequency intracranial decoding above chance [102], word-pair classification from direct cortical recordings [84], non-invasive phrase decoding during imagination stages [29], and recent real-time decoding of attempted and imagined-speech in implanted users [68]. Counter-evidence includes weak acoustic-envelope recoverability under covert conditions [10], declines in decoding performance in earlier continuous speech pipelines [52], and systematic reviews emphasizing heterogeneity in tasks, compliance checks, and preprocessing choices [117]. Methodologically, limited paired neural and speech data for supervision and temporal misalignment persist as practical bottlenecks [20, 123, 49, 141]. Overall, current studies indicate feasibility for proof-of-concept tasks but suggest that clinical-grade reliability remains to be established.
@@
Across multiple datasets, an SNR gradient is observed from overt to mimed (silently attempted without vocalization) to imagined speech, with overt conditions typically exhibiting stronger and more spatially extensive high-gamma (BHA) activity than imagined conditions [102, 10, 114]. This gradient both motivates and constrains imagined-speech decoding: positive feasibility evidence shows above-chance decoding of covert and imagined speech in constrained word and phrase paradigms using low-frequency and CFC features [102, 84, 29], with recent implanted-participant work reporting real-time decoding of attempted and imagined speech [68]. At the same time, negative findings indicate that covert signals encode acoustic detail less reliably and exhibit substantial variability across participants and tasks [10, 52, 117], suggesting that robust large-vocabulary communication is not yet achievable. Crucially, the SNR gradient guides the choice of intermediate representations for decoding. Overt speech-related motor signals are stronger and more directly linked to acoustic features, supporting decoding strategies that exploit high-frequency articulatory correlates [83, 17]. In contrast, imagined speech exhibits weaker, more spatially diffuse activity and poorer discriminability of articulatory representations in BHA; instead, it more reliably engages perceptual (phonetic and vocalic) representations indexed by low-frequency bands and CFC [102, 77, 30], consistent with the flexible abstraction hypothesis [26]. Given the weaker SNR and typical data scarcity in implanted settings, this motivates targeting low-frequency and CFC features [102, 77] and decoding through explicit articulatory–acoustic intermediate parameters (e.g., pitch, formant frequencies) [8] rather than direct spectrograms or latent vectors, which can reduce sample complexity and improve trainability with limited neural data [20, 4]. Thus, imagined-speech variability demands adaptation toward perceptual and low-frequency structure [102, 26, 30], while overt speech can exploit higher-SNR, high-frequency articulatory signals [83, 17].
@@ 2.4 Temporal Dynamics and Context Integration in Speech Representation
The critical role of temporally recurrent connections in cortical regions for generating context-dependent phonological representations is reflected in both computational models and empirical neural data. This principle is computationally realized in the Long Short-Term Memory (LSTM) network introduced by Hochreiter and Schmidhuber in 1997 [54], which, through its memory cells featuring constant error carousels and multiplicative gate units, is designed to control information flow. This architecture enables the storage of context-dependent information over long intervals, thereby facilitating the production of dynamic, context-sensitive phonological representations. Building on this, Graves et al. [41] demonstrated in 2006 that bidirectional LSTM (BiLSTM) networks, trained with Connectionist Temporal Classification (CTC) loss, can implicitly model inter-label dependencies. This is achieved by predicting labels that frequently co-occur in connected patterns, reflecting how cortical recurrent connections may integrate temporal context to shape phonological encoding. The role of recurrent dynamics in speech processing is empirically reinforced by recent studies. Luo et al. (2022) [81] showed that deep learning models like BiLSTMs map temporally structured neural features to sequence-based speech representations, a process underpinned by recurrent cortical connections that enable context-dependent integration [94, 127]. Corroborating this, Metzger et al. (2023) [90] provided evidence that the sensorimotor cortex (SMC) maintains temporally recurrent articulatory representations. These representations are somatotopically organized by place of articulation (POA) and are actively involved in context-dependent phonological encoding.
@@ 3.1 Intracranial Electrode Families, Representative Suppliers, and Development Trajectories
The trajectory of intracranial electrodes in language BCI reflects a progression from clinically available mapping hardware toward purpose-optimized neuroprosthetic interfaces. Early proof-of-concept decoding studies leveraged electrodes originally designed for epilepsy monitoring, beginning with macro-ECoG grids and strips, and later incorporating SEEG depth electrodes, thereby capitalizing on established neurosurgical workflows and clinical monitoring infrastructure [72, 71, 52]. This foundational work demonstrated that cortical field potentials, especially in the high-gamma band, contain rich speech information, and that intracranial recordings are capable of supporting continuous or near-continuous decoding in controlled settings.
The next phase focused on improving spatial resolution, task complexity, and decoding fidelity. High-density cortical recordings and improved modeling pipelines enabled stronger articulation-level decoding and speech synthesis, while broader datasets clarified how electrode placement and coverage heterogeneity shape performance [5, 82, 93, 81, 34]. In parallel, SEEG expanded the accessible anatomical search space beyond cortical surface coverage, supporting distributed language-network sampling and more flexible trajectory planning under clinical constraints [20, 134, 101]. Recent work represents a third phase in which electrode hardware is increasingly evaluated by chronic utility rather than acute signal quality alone. Intracortical MEAs enabled high-rate and high-fidelity speech neuroprosthesis demonstrations [130, 13, 132], while chronic ECoG systems demonstrated stronger home-use feasibility in selected users with ALS [80, 4]. This shift reframes the notion of the “best electrode” as a use-case-dependent decision – one that balances coverage, resolution, packaging, calibration burden, and care logistics, rather than simply maximizing any single signal metric [116, 112, 135].
@@
Intracortical MEAs have also shown significant advancements in capturing speech-related neural activity. Willett et al. (2023) [130] used high-resolution spiking activity recordings from intracortical arrays to achieve a 9.1% WER on a 50-word vocabulary and 23.8% on a 125,000-word vocabulary, with decoding speeds of 62 words per minute. They further identified that the ventral premotor cortex (area 6v) contains a dense, intermixed representation of speech articulators within a small (3.2 × 3.2 mm²) region, with phoneme articulation details preserved even years after paralysis. Notably, decoding accuracy improved log-linearly with electrode count, suggesting that denser arrays could further enhance performance. Similarly, Wood et al. (2025) [132] implanted 256 microelectrodes in the precentral gyrus, capturing neural activity that drove a voice synthesizer to generate speech resembling the participant’s pre-ALS voice with  25 ms latency, including paralinguistic features like intonation and emphasis for modulating speech style (e.g., statements vs. questions). For ECoG, Luo et al. (2022) [81] highlighted innovations such as 32 recording contacts per electrode, robotic insertion techniques to minimize tissue trauma, and flexible array designs, which address traditional limitations in spatial density and coverage. These advances, combined with deep learning methods mapping ECoG high gamma activity to vocal tract trajectories, support improved neural capture for speech decoding. While long-term ECoG stability for speech remains under investigation, motor BCI research has demonstrated reliable chronic decoding, and safety in individuals with late-stage ALS has been reported.
These emerging technologies, including MoS2-based active arrays with ultrahigh spatiotemporal resolution, high-channel-count intracortical microelectrodes, and advanced ECoG configurations, overcome traditional constraints by enhancing resolution, biocompatibility, and coverage. They enable the capture of fine-grained speech-related neural activity, from spiking dynamics to paralinguistic features, and hold significant potential for advancing the clinical translation of language BCIs.